OpenStudy (anonymous):

can someone show me how to find the # of terms required to approximate ln 2 with an error less than 10^(-6)?

OpenStudy (anonymous):

ln 2 is given as \[\sum_{k=1}^{\infty} \frac{ (-1)^{k+1} }{ k }\]

OpenStudy (anonymous):

the magnitude of the remainder is given as,\[\left| Rn \right|\le a _{n+1}\]

OpenStudy (anonymous):

oh nvm, I got it